OpenStudy (anonymous):

The formula for the area of a triangle is A = ½bh where b is the base and h is the height. What is the area of the triangle shown here? right triangle A. 72 in.2 B. 90 in.2 C. 180 in.2 D. 360 in.2

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

Could you post the picture of the triangle?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

it wont let me

OpenStudy (shamil98):

A = 1/2 bh A = (1/2) (15*12) A = (15*12)/2 Can you simplify this?

OpenStudy (anonymous):

wat u mean

OpenStudy (anonymous):

its 180

OpenStudy (shamil98):

15 x 12 = 180 correct 180/2 = your answer

OpenStudy (anonymous):

thank u im bout to post another 1

OpenStudy (anonymous):

Denise has a collection of quarters and dimes. q represents the number of quarters and d represents the number of dimes Denise has in her collection. Which expression describes the amount of money she has in quarters and dimes? A. 0.35dq B. + C. 0.25 + q + 0.1d D. 0.25q + 0.1d

OpenStudy (haseeb96):

D is correct answer

OpenStudy (haseeb96):

because quarter is 1\4 = 0.25 and dimes .1

OpenStudy (anonymous):

ummm witch one ABCD